OpenStudy (anonymous):

Solve the limit as h approaches 0. LIMIT BELOW!

OpenStudy (anonymous):

\[\lim_{h \rightarrow 0}\frac{ 7^{h}-1 }{ h }\]

OpenStudy (anonymous):

I found out the answer if supposed to be like log(7) or ln(7) according to WolframAlpha. Cant see how to get there really. I thought maybe i'd multiply by ln/ln , and get \[\frac{ h-\ln 1 }{ \ln h }\] but that doesn't work because you can't take the natural log of 0 on the bottom still

OpenStudy (anonymous):

this is the derivative of \(7^x\) at \(x=0\) the derivative of \(7^x\) is \(7^x\ln(7)\) so at \(x=0\) you get \(7^0\ln(7)=\ln(7)\)

OpenStudy (anonymous):

i don't think there is a way to do this directly if you have not gotten to derivatives yet. you can do it numerically by trying smaller and smaller values of \(h\) and approximating it with a decimal

OpenStudy (anonymous):

I haven't really learned derivatives yet.. i just know i was suposed to transform that limit so that I could solve it

OpenStudy (anonymous):

you could also use l'hopital, but that presupposes knowing derivatives

OpenStudy (anonymous):

oh ok. well thanks

OpenStudy (anonymous):

you might try rewriting as \[\lim_{n\to \infty}\frac{7^{\frac{1}{n}}-1}{\frac{1}{n}}\] but i am not sure what that is going to buy you.

OpenStudy (anonymous):

well thanks anyway. I guess i can just ask my teacher about it tomorrow